OpenStudy (anonymous):

f(x)=4x+1 and g(x)=x^2-5 , find (f o g)(4) how do I do this??

OpenStudy (agent47):

f(g(x))=f(x^2-5)=4(x^2-5)+1

OpenStudy (agent47):

f(g(4))=4(16-5)+1=4*11+1=45

OpenStudy (jamesj):

@Agent47, of course your answer is correct, but as a general policy, we prefer that you lead people to the answer rather than laying it all out at once.

OpenStudy (agent47):

@JamesJ as far as I can tell, this place hasn't changed a bit, and almost nobody ever leads people to the answer rather than laying it all out at once.

OpenStudy (jamesj):

Well, you can be part of the difference then!

OpenStudy (lgbasallote):

@SantanaG another method is to find g(4) first. what do you think is g(4)?

OpenStudy (anonymous):

g(4)= 4^2-5 =11

OpenStudy (lgbasallote):

correct. now find f(11)

OpenStudy (anonymous):

4(11)+1=45

OpenStudy (lgbasallote):

that's the other method :)

OpenStudy (anonymous):

oh wow that was easy!! thank you @lgbasallote

OpenStudy (lgbasallote):

hehe lol. just make sure you dont get confused with using that :) it can sometimes get tricky